OpenStudy (anonymous):

A Daredevil Jumps Off A Tall Building. He Falls Freely For Several Seconds Before Releasing His Parachute . His Height (h) In Meters , T Seconds After Jumping Can Be Modeled By : h=-4.9t^2+t+360 before he releases his parachute and h=-4t+142 After He Releases His Parachute . How Long After Jumping Did He Release His Parachute ? HELPP ME PLEASE !!!!

OpenStudy (anonymous):

PLEASE HELP ! THIS IS A BIG QUESTION WORTH LOSTS OF POINTS

OpenStudy (noelgreco):

Although it doesn't say so, the second equation suggests he's landed. Solve for t.

OpenStudy (anonymous):

Ok So -4.9t + t + 360=4t +142

OpenStudy (noelgreco):

Actually, I put my glasses on an d saw that all you have to do is assume the h valuues are the same, equate the to equations, and solve for t

OpenStudy (anonymous):

im confused

OpenStudy (irishboy123):

you have \( h_1=-4.9t^2+t+360\) \(h_1(0) = 360\) he jumped off a 360m tall building and then you have \(h_2=-4t+142\) \(h_2(0) = 142\) he pulled the chute at 142 m above ground so go back to \(h_1(t)\) and work out how long it took him to fall from 360m to 142m you have a quadratic to solve